Operation Summary for June 13, 2011 - June 15, 2011 |
|
Monday June 13 The day shift began with Operations monitoring the Tevatron (TEV) store 8811 (172.71E30) and the Recycler stash (152.1E10), with the Antiproton Source (Pbar) stacking (18.7E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations reported a problem with MECAR pulling the MI beam permit; MI personnel investigated and resolved the problem. Cryo system technicians worked on the TEV sectorB2 cryo system. MI personnel conducted a coalescing check. During the evening shift, Operations terminated beam to T-978 at 6:02 PM. Operations terminated store 8811 at 6:54 PM, put the TEV into HEP shot setup, and established store 8812 at 8:02 PM with an average initial luminosity (AIL) of 398.76E30. Tuesday June 14 The midnight shift began with Operations monitoring store 8812 (183.312E30) and the stash (120.7E10), with Pbar stacking (32.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 2:59 AM, operators reported that MECAR caused MI to abort during a transfer of antiprotons from Pbar to the Recycler. Operations couldn’t find a reason for the abort; everything reset okay. Operations established beam to T-978 at 4:35 AM. At 6:39 AM, a TEV RF station (TRF3) tripped off and caused a loss of luminosity: ~8E30. MECAR caused another MI abort at 7:11 AM; everything reset. Operators reported some short downtimes due to a Linac Klystron RF station (KRF1), LRF1, and MIRF15. During the day shift, Operations terminated 8812 at 9:15 AM. Kicker personnel began working on the sweeper magnet at AP0. TEV RF personnel began work on TRF3. Operators prepared CDF and D0 for a collision hall access at 9:28 AM; both experiments came out of access at 11:10 AM. Operations put the TEV through a dry squeeze at 11:39 AM. Operators reported that TRF3 was back online. Operations put the TEV into shot setup at 2:46 PM and established store 8816 at 3:46 PM with and AIL of 424.56E30. During the evening shift, Operations disabled beam to T-978 at 6 PM. Operators reported using the Tevatron Electron Lens (TEL2) for a Hollow Beam study during the shift. Wednesday June 15 The midnight shift began with Operations monitoring store 8816 (125.57E30) and the stash (264.1E10), with Pbar stacking (13.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:20 AM. At 4:54 AM, Operations reported that store 8816 aborted due to a lightning strike that caused a Fast Bypass interlock loop trip. Operators investigated and found a volts-to-ground indication. Operators put the TEV through a dry squeeze and a wet squeeze. Operations put the TEV into shot setup at 6:56 AM and established store 8818 at 7:56 AM with an AIL of 376.02E30. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct a beam cooling test and Linac personnel to conduct a beam scan. |
|
More Information Comments and Suggestions |
| last modified 06/15/2011 email Fermilab |
| Security, Privacy, Legal | |